OpenStudy (anonymous):

how to answer this question? a) show that the composite function of two one-to-one functions is also one-to-one function? how to answer this question? a) show that the composite function of two one-to-one functions is also one-to-one function? @Mathematics

jhonyy9 (jhonyy9):

so i think that is the same like when we have one square perfect like (x+1) squared and calculating we get x2+2x+1

OpenStudy (slaaibak):

no. (x + 1) squared is NOT one-to-one. It should go like this: suppose you have two one-to-one functions: f(x) = x + 1 g(x) = 2x + 3 f(g(x)) = (2x + 3) + 1 = 2x + 4 -----> one-to-one g(f(x)) = 2(x + 1) + 3 = 2x + 5 -----> one-to-one this shows the composite function of two one-to-one functions is also a one-to-one function.

OpenStudy (anonymous):

thanx alot guys ! its really helpful :D